# Reliable uncertainty with cheaper neural network ensembles: a case study in industrial parts classification

### v1

@@ 4.2 Data
We use the Synthetic Industrial Parts (SIP-17) dataset (Zhu et al., 2023, cases 1 and 2), a dataset containing images of various industrial parts. The ID set consists of five industrial parts labeled “Airgun,” “Electricity12,” “Hammer,” “Hook,” and “Plug,” while the OOD set consists of three industrial parts labeled “Fork1,” “Fork2,” and “Fork3” (Figure 4).
The ID set is split in a training (70%), validation (15%), and test set (15%). The OOD set is entirely used as test set (i.e., 100%). The dataset is normalized to the [0,1] range, resized to 32$\times$32 pixels, and converted to grayscale. During training, data augmentation is applied with random horizontal and vertical flips.
@@
4.3 Models
Throughout the experiments, the base model is a LeNet-5 with $80,865$ parameters in the standard configuration. All models except the snapshot ensemble are trained using the Adam optimizer with an initial learning rate of $0.001$; the snapshot ensemble uses the custom cyclic learning rate schedule with a maximum learning rate of $0.001$. The single model, deep ensemble, and snapshot ensemble are trained for 200 epochs. Following Nado et al. (2021), the batch ensemble and MIMO ensemble are trained for 25% longer (i.e., 250 epochs) because they take longer to converge. All models are trained with a batch size of 512. The batch ensemble uses random signed vectors for the fast weights. The MIMO ensemble uses an input repetition probability of 0.8. The deep, snapshot, and batch ensemble have 10 ensemble members. In contrast, the MIMO ensemble only has 4 members as the capacity of the individual members is otherwise too limited. For completeness, Appendix A compares results between MIMO ensembles comprising 4 and 10 members, while Appendix B shows results when all ensembles consist of 4 members.
@@ 4.3 Models
The models are implemented in TensorFlow using the Uncertainty Baselines repository (Nado et al., 2021) and are trained on an NVIDIA RTX A5000 card. The results are averaged over 5 independent runs with random seeds for a total runtime of 8 hours.
4.4 Evaluation metric: Diversity Quality
@@
To gain a deeper understanding of the ensembling techniques, we assess the diversity among ensemble members. The diversity metric measures the fraction of test data points on which predictions of ensemble members disagree (Fort et al., 2019). This metric is 0 when two members make identical predictions and 1 when predictions differ on every example in the test set. As the base model for diversity computation, we average the output distributions over the members and determine the resulting predicted label. The diversity score is related to the model uncertainty in section 5.2 and allows to assign a score to each member.
@@
We propose a novel Diversity Quality ($DQ_{1}$) score to represent the diversity on both the ID and OOD set in one single metric, made available in the Python package reject (Thuy and Benoit, 2024). The ID diversity (IDD) should be as close as possible to 0.0 and the OOD diversity (OODD) as close as possible to 1.0. The $DQ_{1}$-score is the harmonic mean of (1 - IDD) and OODD, displayed in Equation 6. The harmonic mean assigns equal weights to both IDD and OODD and is most appropriate to handle ratios (Agrrawal et al., 2010). It has a minimum value of 0.0 and a maximum value of 1.0. The new evaluation score has the advantage of being close to 0.0 when either the IDD or OODD is poor, while at the same time, the score is close to 1.0 only when both the IDD and OODD are strong.
@@ 5.1 In-distribution performance
We evaluate the classification accuracy and the negative log-likelihood (NLL) for the ID predictions. The NLL is commonly used to measure the quality of the model’s uncertainty (Ovadia et al., 2019); lower scores are better. For ensembles, predictions are averaged over the members before metrics are calculated. We expect the ensemble methods to perform better than the single NN. Table 2 shows the results.
@@
Single 96.61 $\pm$ 0.18 0.1121 $\pm$ 0.0039
Deep E ($M=10$) 98.12 $\pm$ 0.13 0.0695 $\pm$ 0.0024
Snapshot E ($M=10$) 94.69 $\pm$ 0.47 0.1774 $\pm$ 0.0096
Batch E ($M=10$) 98.79 $\pm$ 0.11 0.0499 $\pm$ 0.0029
MIMO E ($M=4$) 95.96 $\pm$ 0.44 0.1168 $\pm$ 0.0092
In terms of accuracy, the deep ensemble demonstrates a significant performance advantage over the single model. Among the efficient ensembles, the snapshot ensemble falls short, exhibiting substantially poorer performance compared to the single model. Conversely, the batch ensemble displays robust performance and even surpasses the deep ensemble. The MIMO ensemble achieves a slightly lower accuracy than the single model. Similar conclusions hold for the NLL metric, where once again, the batch ensemble outperforms the deep ensemble. A lower NLL score indicates a closer alignment of predicted probabilities with the true labels. Overall, the batch ensemble presents highly promising results on the ID set, surpassing the benchmark set by the deep ensemble. Subsequent sections will delve into the OOD behavior of the models.
@@ 5.2 Uncertainty density
(b) Data uncertainty
(c) Model uncertainty
Figure 5: Densities for the data, model, and total uncertainty on the ID and OOD set
Figure 5 illustrates the density of total, data, and model uncertainty. The results are obtained by evaluating the models on the ID and OOD set, and calculating the uncertainty in each prediction. Note that total uncertainty is the sum of data and model uncertainty and has a maximum value of $\log_{2}(5)=2.32$. We do not use the NLL here because this requires ground truth labels unavailable for the OOD set. On the ID data, we would like to see low total uncertainty. Conversely, on OOD data, we expect the models to exhibit higher model uncertainty and thus total uncertainty, reflecting an awareness that they know what they do not know. In a highly idealized scenario, the mode of the total uncertainty lies around zero for the ID set and approaches the maximum value of 2.32 for the OOD set.
The single model displays low uncertainty on the ID set but only slightly higher uncertainty on the OOD set. Notably, this uncertainty stems solely from data uncertainty with a mode around 0.2, as the single model lacks the ability to capture model uncertainty. In contrast, the deep ensemble aligns with expectations, demonstrating low uncertainty on the ID set and substantial model uncertainty on the OOD set, resulting in higher total uncertainty with a mode around value 1.0.
Among the efficient ensembles, we observe significant variations in uncertainty behavior. The snapshot ensemble exhibits surprisingly high total uncertainty on the ID set, resulting from high data uncertainty which is undesirable. Furthermore, it has higher total uncertainty on the OOD set with a mode around 1.0 but this primarily stems from data uncertainty, not model uncertainty. The batch ensemble performs well, mirroring the behavior of the deep ensemble with low uncertainty on the ID set and increased model uncertainty on the OOD set with a mode around 1.2. Lastly, the MIMO ensemble indicates low uncertainty on the ID set but falls short on the OOD set. Here, the OOD set induces higher data uncertainty while model uncertainty only experiences a slight increase. Consequently, the total uncertainty on the OOD set has a mode around value 0.75, which is limited compared to other ensemble techniques.
Overall, the batch ensemble once again stands out as the most promising, exhibiting similar uncertainty behavior to the Deep ensemble. Conversely, the snapshot and MIMO ensemble do not show substantially increased model uncertainty on the OOD set.
@@ 5.3 Classification with rejection
In practical scenarios, it is crucial for a model to avoid overly confident yet incorrect predictions. One method to address this involves passing predictions with high uncertainty to human experts for labeling, known as classification with rejection (Barandas et al., 2022). To assess the ensembles’ uncertainty estimates, accuracy is computed only for observations with a total uncertainty value below a user-defined threshold, referred to as non-rejected accuracy (NRA).
We evaluate the models on the combination of the ID and OOD test sets. A model proficient in identifying OOD data exhibits an NRA curve closer to the top-left corner. Essentially, as the NRA curve rises earlier, the model demonstrates higher uncertainty on observations that in fact prove to be OOD data. Consequently, it prioritizes rejecting OOD observations, then misclassified ID observations, and finally correctly classified ID observations at lower uncertainty levels. Figure 6 displays the NRA for all models based on their total uncertainty.
@@
The single model exhibits the weakest performance, characterized by a slow increase in NRA, reaching only 40% at low uncertainty thresholds. This finding aligns with the low uncertainty observed on the OOD set in Figure 5. In contrast, the deep ensemble demonstrates significantly better performance, showing a much earlier and sharper increase in NRA, reaching 80% at low uncertainty thresholds.
The snapshot ensemble slightly underperforms the deep ensemble until a threshold of 0.25; it is only for smaller uncertainty thresholds that snapshot ensemble has a rapid improvement in NRA. The batch ensemble demonstrates strong performance, outperforming the deep ensemble across the entire range of thresholds. Conversely, the MIMO ensemble disappoints with a slowly increasing NRA, placing it between the single model and deep ensemble.
In summary, the single model produces confident incorrect predictions, as indicated by the low NRA curve. The deep ensemble is substantially better at identifying and rejecting OOD observations, demonstrated by its early and steep increase in NRA. Both the snapshot and batch ensemble perform well and match or even improve on the deep ensemble. Unfortunately, the MIMO ensemble falls short of achieving equally strong NRA values.
@@ 5.4 Diversity analysis
On the ID set (Figure 6(a)), the MIMO ensemble performs best and has the lowest diversity, followed by the batch ensemble. Both the snapshot and deep ensembles demonstrate slightly higher diversity scores. Notably, the snapshot ensemble displays a skewed diversity distribution, with several members exhibiting significantly higher diversity than the majority. This behavior stems from the initial snapshots being highly diverse, gradually converging as training progresses.
Moving to the OOD set (Figure 6(b)), the diversity scores are indeed orders of magnitude higher. Here, the MIMO ensemble still has the lowest diversity, which is undesirable as the diversity should be as high as possible on the OOD set. The batch ensemble has the highest diversity and performs strongest, closely followed by the deep ensemble. Again, the snapshot ensemble has generally low diversity, with early snapshots being very diverse.
@@
Figure 8 shows the $DQ_{1}$-scores for each ensemble and its members. The batch ensemble stands out with the highest $DQ_{1}$-scores, attributed to its relatively low IDD and high OODD. The deep ensemble follows closely behind with marginally lower scores than the batch ensemble because of slightly higher IDD and lower OODD. Furthermore, the snapshot ensemble again exhibits a skewed distribution of scores due to its unique training dynamics. The MIMO ensemble ranks lowest in $DQ_{1}$-scores because of its undesirable low OODD, although it has advantageous low IDD. This lack of diversity is reflected in its overall poor $DQ_{1}$-scores. Note that the single NN has a minimum score of zero because it only has one member, thereby lacking diversity altogether. Overall, the $DQ_{1}$-scores comprehensively evaluate the ensembles’ diversity performance.
@@ 5.5 Computational cost
Figure 9 depicts the computational costs associated with all ensembling techniques. The values are normalized relative to the single model to facilitate easier comparisons.
@@
In Figure 8(a), the training cost is presented. The training time for the deep ensemble with $M=10$ members is 10 times longer than that of the single model, as it involves training 10 NNs independently. The training time for the snapshot ensemble is equivalent to that of a single model, and the number of ensemble members does not impact the training duration. Both the batch ensemble and MIMO ensemble exhibit slightly slower training times compared to the single model, with a factor of 1.5x and 1.25x, respectively. It is important to note that the batch ensemble and MIMO ensemble are trained for 25% more epochs due to a slower learning convergence.
In Figure 8(b), evaluation times are presented. The deep and snapshot ensembles incur 10 times higher evaluation costs than the single NN because they require $M=10$ independent forward passes. Conversely, the batch ensemble and MIMO ensemble are only around 20% slower in evaluation.
Figure 8(c) illustrates the number of parameters stored. Memory consumption for both the deep and snapshot ensembles is 10 times higher, given the weights storage for $M=10$ individual NNs. The batch ensemble has approximately 10% more parameters due to the inclusion of additional fast weights. In the MIMO model, the extra parameters in the input and output layers are negligible.
In summary, the batch ensemble and MIMO ensemble incur only slight additional costs compared to the single model, making them orders of magnitude more cost-effective than the deep ensemble.

### v2

@@ 5.1 Data
We use the Synthetic Industrial Parts (SIP-17) dataset (Zhu et al., 2023) containing synthetic images of 17 industrial parts. Although the images are synthetic, SIP-17 allows for simulating OOD scenarios frequently encountered in real-life OR scenarios.
Large enterprises have inventories reaching thousands of items (Ernst and Cohen, 1990) with operational and maintenance support expenses often representing over 60% of the total costs (Hu et al., 2015). As such, it is infeasible for human operators to manually identify each part and its appropriate inventory strategy. To this end, identification is often based on barcodes but is prone to wrong labeling, parts may be too small, or functional surfaces do not allow the attachment of a code label. Using explicit codes is especially challenging in remanufacturing processes of disassembled products, where barcodes are likely dirty or damaged, limiting the automation, reliability, and quality of the processes. Industrial parts classification offers a solution by identifying parts based on their visual features, maintaining the efficient automation of logistic processes such as part supply and sorted storing. Depending on the specific setup, a camera can be mounted above a conveyor belt or the camera of a tablet can be used.
The images of the 17 industrial parts are divided into six use cases. The images were generated by rendering 3D CAD models from varying camera angles and under diverse lighting conditions.
5.2 Experimental setup
Figure 4: Experimental setup
Figure 4 outlines the experimental setup, consisting of an ID and OOD test set. The models are trained on synthetic images of case 1. First, we use a synthetic test set of the same distribution, i.e., the ID test set. Second, we evaluate on synthetic images of the objects in cases 2–4 not observed during training, i.e., the OOD test set. This set simulates situations such as encountering newly developed parts introduced after the training phase or parts managed by different departments. It is imperative that the classifier can identify these novel situations and signal a warning, prompting human intervention, rather than making incorrect predictions.
@@ 5.2 Experimental setup
Table 3: Number of images per dataset
Set Train Validation Test
ID 5100 900 1500
OOD 3000
Figure 5 shows an example image for each object type in the two sets. The ID set consists of 5 industrial parts while the OOD set contains 10 industrial parts; the dataset is balanced. Note that we do not employ cases 5 and 6 as OOD data because these cases contain images of assembled parts, e.g. the front and rear view of a wheel with and without a wheel nut. As such, these images are highly similar which would cause the wheel object to be overrepresented in the OOD set. Table 3 displays the dataset sizes; an overview of the dataset size per object type is available in Appendix A. The ID set available for training is split into a training (85%) and validation (15%) set; the OOD set is entirely used as test sets (i.e., 100%). We investigate the predictive performance and uncertainty quality of all models, along with their computational expense.
5.3 Implementation
Images are resized to 32$\times$32 pixels in grayscale and the dataset is normalized to the [0,1] range. During training, data augmentation is applied with random horizontal and vertical flips.
Throughout the experiments, the base model is a LeNet-5 with $80,865$ parameters in the standard configuration. LeNet-5 is a popular CNN architecture which has been shown to work well on smaller input images (Chang, 2023). Given that for deep and batch ensemble larger ensemble sizes generally translate to better performance, we use $M=8$ members and provide additional results for $M=4$ in Appendix C. For snapshot and MIMO ensemble, selecting the ensemble size $M$ is more complex as using too many or too few members might negatively impact the predictive or uncertainty performance. Following the authors’ guidelines, we tune snapshot ensembles with 4, 6, and 8 members and select the best-performing configuration based on the validation set. Similarly, we tune MIMO ensembles with 3 and 4 members and select the best-performing configuration. A detailed overview of the hyperparameters is available in Appendix B, together with the validation performance of each configuration.
@@ 5.3 Implementation
The models are implemented using the Uncertainty Baselines repository (Nado et al., 2021) and are trained on an NVIDIA RTX A5000 GPU. The results are averaged over 5 independent runs with random seeds for a total runtime of 40 hours.
5.4 Evaluation metrics
In addition to the classification accuracy $\uparrow$ (the arrow indicates which direction is better), we examine the predictive uncertainty behavior. Evaluating the predictive uncertainties is challenging as the “ground truth” uncertainty estimates are unavailable. In this work, we use calculate the calibration metric negative log-likelihood (NLL) on the ID set. However, this cannot be calculated on the OOD set because there is no ground truth label and the predictions are incorrect by definition. Uncertainty quantification can be evaluated using classification with rejection and the novel Diversity Quality metric.
Negative log-likelihood (NLL) $\downarrow$ NLL is a proper scoring rule and a popular metric for evaluating predictive uncertainty (Ovadia et al., 2019). For $N$ observations and $K$ output classes, this gives:
$\textrm{NLL}=\frac{1}{N}\sum_{n=1}^{N}\sum_{k=1}^{K}\mathbbm{1}(y_{n}=k)\log p(y=k\mid\mathbf{x}_{n},\ \bm{\theta}).$ (6)
A lower NLL score indicates a closer alignment of predicted probabilities with the true labels.
[…]
@@ 5.4 Evaluation metrics
Diversity Quality metric $\uparrow$ To gain a deeper understanding of the ensembling techniques, we assess the diversity among ensemble members. The diversity metric measures the fraction of test data points on which predictions of ensemble members disagree (Fort et al., 2019). This metric is 0 when two members make identical predictions and 1 when predictions differ on every example in the test set. As the base model for diversity computation, we average the output distributions over the members and determine the resulting predicted label. The diversity score is related to the epistemic uncertainty in Section 6.2 and allows assigning a score to each member.
@@
We propose a novel Diversity Quality ($DQ_{1}$) score to represent the diversity performance on both the ID and OOD set in one single metric, made available in the Python package reject (Thuy and Benoit, 2024). The ID diversity (IDD) should be as close as possible to 0.0 and the OOD diversity (OODD) as close as possible to 1.0. The $DQ_{1}$-score is the harmonic mean of (1 - IDD) and OODD, displayed in Equation 7. The harmonic mean assigns equal weights to both IDD and OODD and is most appropriate to handle ratios (Agrrawal et al., 2010). It has a minimum value of 0.0 and a maximum value of 1.0. The new evaluation score has the advantage of being close to 0.0 when either the IDD or OODD is poor, while at the same time, the score is close to 1.0 only when both the IDD and OODD are strong.
@@
The $DQ_{1}$-score is beneficial for practitioners because it simplifies the evaluation of diversity performance. This point is important as the underlying IDD and OODD metrics are difficult to assess separately because they often exhibit a trade-off. Additionally, the $DQ_{1}$-score allows practitioners to consistently prioritize either IDD or OODD by adjusting the $\beta$ value. For instance, in medical applications where incorrect predictions can have severe consequences, practitioners may prioritize achieving high OODD to minimize the risk of incorrect treatments in unseen situations.
@@ 6 Results and Discussion
This section examines the results for both the ID and OOD sets, focusing on predictive performance, uncertainty behavior, Diversity Quality, and computational cost. Finally, we address the trade-offs among these aspects to identify the most suitable NN ensemble for practical applications. We work with a snapshot ensemble of $M=8$ members and a MIMO ensemble of $M=3$ members because this yields the lowest validation NLL for their ensemble type. Please refer to Appendix B for an overview of the validation performance.
@@ 6.1 In-distribution performance
We evaluate the classification accuracy and NLL for the ID predictions. We expect the ensemble methods to perform better than the single NN. Table 4 shows the results.
@@
Single 97.72 $\pm$ 0.14 0.1047 $\pm$ 0.0089
Deep E ($M=8$) 98.53 $\pm$ 0.09 0.0440 $\pm$ 0.0011
Snapshot E ($M=8$) 97.15 $\pm$ 0.10 0.1048 $\pm$ 0.0031
Batch E ($M=8$) 98.65 $\pm$ 0.05 0.0504 $\pm$ 0.0035
MIMO E ($M=3$) 96.63 $\pm$ 0.30 0.1057 $\pm$ 0.0061
In terms of accuracy, the deep ensemble outperforms the single neural network (NN) significantly. Among the efficient ensembles, both the snapshot ensemble and MIMO ensemble underperform, delivering worse results compared to the single NN. In contrast, the batch ensemble demonstrates strong performance, matching that of the deep ensemble. Similar trends are observed for the NLL metric, where the batch ensemble again matches the deep ensemble, while the snapshot and MIMO ensembles perform similarly to the single NN. Overall, the batch ensemble shows highly promising results on the ID set, equaling the benchmark set by the deep ensemble. The following sections will explore the models’ behavior on OOD data.
@@ 6.2 Uncertainty density
(b) Aleatoric uncertainty
(c) Epistemic uncertainty
Figure 6: Densities for the aleatoric, epistemic, and total uncertainty on the ID and OOD set
Figure 6 illustrates the density of total, aleatoric, and epistemic uncertainty. The results are obtained by evaluating the models on the ID and OOD sets, and calculating the uncertainty in each prediction. Note that total uncertainty is the sum of aleatoric and epistemic uncertainty and has a maximum value of $\log_{2}(5)=2.32$. We do not use the NLL here because this requires ground truth labels unavailable for the OOD set. On the ID data, we would like to see low total uncertainty. Conversely, on OOD data, we expect the models to exhibit higher epistemic uncertainty and thus total uncertainty, reflecting an awareness that they know what they do not know. In a highly idealized scenario, the mode of the total uncertainty lies around zero for the ID set and approaches the maximum value of 2.32 for the OOD set.
The single NN displays low uncertainty on the ID set but only slightly higher uncertainty on the OOD set. Notably, this uncertainty stems solely from aleatoric uncertainty with a mode around 0.0, as the single NN lacks the ability to capture epistemic uncertainty. In contrast, the deep ensemble aligns with expectations, demonstrating low uncertainty on the ID set and substantial epistemic uncertainty on the OOD set, resulting in higher total uncertainty with a mode around value 0.95.
Among the efficient ensembles, we observe variations in uncertainty behavior. The snapshot ensemble exhibits moderately high total uncertainty on the ID set. Furthermore, total uncertainty increases on the OOD set with a mode around 0.75 but this primarily stems from aleatoric uncertainty, not epistemic uncertainty which is undesirable. The batch ensemble performs well, mirroring the behavior of the deep ensemble with low uncertainty on the ID set and increased epistemic uncertainty on the OOD set resulting in a mode of total uncertainty around 1.0. Lastly, the MIMO ensemble exhibits decent uncertainty values on both the ID and OOD sets. Here, the total uncertainty on the OOD set has a mode around value 0.85 but this is mainly driven by higher aleatoric uncertainty while epistemic uncertainty only experiences a slight increase.
Overall, the batch ensemble once again stands out as the most promising, exhibiting similar uncertainty behavior to the deep ensemble. Conversely, the snapshot and MIMO ensemble do not show substantially increased epistemic uncertainty on the OOD set.
@@ 6.3 Classification with rejection
We evaluate the models on the combination of the ID and OOD test sets. A model proficient in identifying OOD data exhibits an NRA curve closer to the top-left corner. Essentially, as the NRA curve rises early and monotonically, the model demonstrates higher uncertainty on observations that in fact prove to be OOD data. Consequently, it prioritizes rejecting incorrectly classified ID observations or OOD observations before rejecting the correctly classified ID observations. Figure 7 displays the NRA for all models based on their total uncertainty.
@@
The single NN exhibits the weakest performance, characterized by a slow increase in NRA, reaching only 45% at low uncertainty thresholds. This finding aligns with the low uncertainty observed on the OOD set in Figure 6. In contrast, the deep ensemble demonstrates significantly better performance, showing a much earlier and sharper increase in NRA, reaching 65% at low uncertainty thresholds.
The snapshot ensemble consistently underperforms the deep ensemble. The batch ensemble demonstrates strong performance, significantly outperforming the deep ensemble across the entire range of thresholds. The MIMO ensemble closely matches the deep ensemble and outperforms it for smaller uncertainty thresholds.
In summary, the single NN produces confident incorrect predictions, as indicated by the low NRA curve. The deep ensemble is substantially better at identifying and rejecting OOD observations, demonstrated by its early and steep increase in NRA. Both batch and MIMO ensemble perform well and match or even improve on the deep ensemble. The snapshot ensemble falls short of achieving equally strong NRA values.
@@ 6.4 Diversity analysis
On the ID set (Figure 8(a)), the deep and batch ensembles perform best and have the lowest diversity, closely followed by the MIMO ensemble. The snapshot ensemble exhibits a skewed diversity distribution, with most members having very low diversity but others having very high diversity. This behavior stems from the initial snapshots being highly diverse and gradually converging as training progresses.
Moving to the OOD set (Figure 8(b)), the diversity scores are indeed orders of magnitude higher. Here, as on the ID set, the snapshot ensemble still has low diversity for most of its members and high diversity for a few early snapshots, which is undesirable as the diversity should be as high as possible on the OOD set. The deep and batch ensemble have the highest diversity and perform strongest, closely followed by the MIMO ensemble.
@@
Figure 9 shows the $DQ_{1}$-scores for each ensemble and its members. The deep and batch ensemble stand out with the highest $DQ_{1}$-scores, attributed to their relatively low IDD and high OODD. Although both ensembles exhibit similar behavior, the deep ensemble has more variation in the diversity scores. The snapshot ensemble ranks lowest in $DQ_{1}$-scores because of its undesirable low OODD, despite having low IDD for most members. This lack of diversity is reflected in its overall poor $DQ_{1}$-scores. Furthermore, the MIMO ensemble slightly underperforms deep and batch ensemble because its IID and OODD are slightly worse. Note that the single NN has a minimum score of zero because it only has one member, thereby lacking diversity altogether. Overall, the $DQ_{1}$-scores comprehensively evaluate the ensembles’ diversity performance.
@@ 6.5 Computational cost
Figure 10 depicts the computational costs associated with all ensembling techniques. The values are normalized relative to the single NN to facilitate easier comparisons.
@@
In Figure 10(a), the training cost is presented. The training time for the deep ensemble with $M=8$ members is 8 times longer than that of the single NN, as it involves training 8 NNs independently. The training time for the snapshot ensemble is equivalent to that of a single NN, and the number of ensemble members does not impact the training duration. Both the batch ensemble and MIMO ensemble exhibit slightly slower training times compared to the single NN, with a factor of 1.85x and 1.50x, respectively. It is important to note that the batch ensemble and MIMO ensemble are trained for 50% more epochs due to slower learning convergence.
In Figure 10(b), evaluation times are presented. The deep and snapshot ensembles incur 8 times higher evaluation costs than the single NN because they require $M=8$ independent forward passes. Conversely, the batch ensemble and MIMO ensemble are only around 50% and 15% slower in evaluation, respectively.
Figure 10(c) illustrates the number of parameters stored. Memory consumption for both the deep and snapshot ensembles is 8 times higher, given the weights storage for $M=8$ individual NNs. The batch ensemble has approximately 10% more parameters due to the inclusion of additional fast weights. In the MIMO model, the extra parameters in the input and output layers are negligible.
In summary, the batch ensemble and MIMO ensemble incur only slight additional costs compared to the single NN, making them orders of magnitude more cost-effective than the deep ensemble.